OpenStudy (anonymous):

Linear programming - how to find the shaded region Graph will be given below.. gimme a moment

OpenStudy (anonymous):

Cake x = 200 g of floor and 25 g of fat cake y = 100 g of floor and 50 g of fat subject to constraints 5 kg of floor(5000g) 1 kg of fat(1000g) Z = x + y 200x + 100y = 5000 - eq1 25x + 50y = 1000 - eq2 eq 1 = x = 25, y = 0 and x = 0, y = 50 eq 2 = x = 40, y = 0 and x = 0, y = 20  where do I shade?